What was the standard cost per unit of product? (Round answer to 2 decimal places, e-g. 2.75. $ Standard cost per unit


Your answer is correct. What were the standard hours allowed for the units produced? 4640 hours Standard hours allowed
tem. In 2020, the company produced 27,900 units. Each unit took several pounds of direct materials ndard hourly rate of $13.00. Normal capacity was 49,700 direct labor hours. During the year, 1.30.300 pounds and 1.6 standard hours of direct labor at a purchased at $0.93 per pound. All materials purchased were used during the year
What was the standard cost per unit of product? (Round answer to 2 decimal places, e-g. 2.75. $ Standard cost per unit
If the 2.75.) Is price variance was $2,606 favorable, what was the standard materials price per pound? (Round answ $ 0.95 Standard materials price per pound
Your answer is correct r was the standard material If the materials quantity variance was $17,765 unfavorable, what was the standard materials quantity per unit? Standard materials quantity per unit pounds
Your answer is correct. What were the standard hours allowed for the units produced? 4640 hours Standard hours allowed
quantity variance was $6,500 unfavorable, what were the actual direct labor hours worked? If the labor 45140 hours Actual hours worked
If the labor price variance was $9,028 favorable, what was the actual rate per hour? (Round answer to 2 decimal places, eg $ 12.80 Actual rate per hour
Your answer is correct. If total budgeted manufacturing overhead was $328,020 at normal capacity, what was the predetermined overhead rate based on direct labor hours? (Round answer to 2 decimal places, e.g. 2.75 Predetermined overhead rate S$
Answer #1
Actual Output = 27900 units
Standard Labour Hours = 1.6 Hrs. per Unit = 27900*1.6 = 44640 hrs.
Standard Labour Rate = $ 13 per hour
Standard Labour Cost = 44640*13 = 580320
Actual Qty. Used = 130300 pounds
Actual Cost per pound = $0.93
Actual Material Cost = 130300*0.93 = 121179
Material Price Variance = 2606 Favourable
Material Qty. Variance = 17765 Adverse
Material Cost Variance = 15159 Adverse
Material Cost Variance = Standard Material Cost (SMC)-Actual Material Cost
-15159 = SMC-121179
SMC = 121179-15159
SMC = 106020
Total Standard Cost = Standard labour Cost+Standard Material Cost
= 580320+106020
= 686340
Standard Cost per unit = Total Standard Cost/Actual Output
= 686340/27900
= 24.6
Material Price Variance = (Standard material price (SMC)-Actual material price)*Actual Qty.
2606 = (SMC-0.93)*130300
SMC = (2606/130300)+0.93
SMC = 0.02+0.93
SMC per pound = $0.95
Material Qty Variance = (Standard material Qty (SMQ)-Actual material Qty)*Standard Material Price
-17765 = (SMQ-130300)*0.95
SMQ = (-17765/0.95)+130300
SMQ = 130300-18700
SMQ = $111,600.00
SMQ per Unit = SMQ/Actual ouput
SMQ per Unit = 111600/27900
SMQ per Unit = 4
Standard Hours allowed for the unit produced = Standard labour hours per unit*Actual output
= 1.6*27900
= 44640
Labour Qty Variance = (Standard Hours Worked -Actual Hours Worked(AHW))*Standard Labour Rate
-6500 = (44640-AHW)*13
AHW = 44640+(6500/13)
AHW = 44640+500
AHW = 45140
Labour Price Variance = (Standard Labour Rate -Actual Labour Rate (ALR))*AHW
9028 = (13-ALR)*45140
ALR = 13-(9028/45140)
ALR = 13-0.20
ALR = 12.80
Total Budgeted Manfacturing Overhead = $328,020
Direct Labour Hour at Normal Capacity = 49700
Predetermined Rate = Total Budgeted Manfacturing Overhead/Direct Labour Hour at Normal Capacity
Predetermined Rate = 328020/49700
Predetermined Rate = 6.6